OpenStudy (judson_rockets):

What is the volume of this triangular prism? A. 432 cm3 B. 38 cm3 C. 216 cm3 D. 864 cm3

OpenStudy (abdullahm):

Volume of a triangular prism = \(\bf \frac{1}{2} \times length \times width \times height\)

OpenStudy (abdullahm):

The length is 8 The width is 12 The height is 18 Plug it in and tell me what you get! :)

OpenStudy (judson_rockets):

1728

OpenStudy (anonymous):

the answer is D

OpenStudy (abdullahm):

@Judson_rockets You forgot to divide by 2. What is 1728/2 = ?

OpenStudy (judson_rockets):

Oh so it is D then

OpenStudy (abdullahm):

That is correct! :)